OpenStudy (anonymous):

Which of the following lists is in order from smallest to largest? , 0.6, 6.2%, , , 0.6, 6.2% 6.2%, 0.6, , 0.6, 6.2%, ,

OpenStudy (igreen):

You're missing some values

OpenStudy (anonymous):

ok |

OpenStudy (anonymous):

or would it

OpenStudy (anonymous):

what

OpenStudy (anonymous):

it c cause .6=60% and 6% is less

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

im not supposed to give answers so i need to know if you understand honestly how to do this

OpenStudy (anonymous):

yes thank you for helping

OpenStudy (anonymous):

no problem ;)

OpenStudy (anonymous):

~Coughs give him a medal @hunter.glover

OpenStudy (anonymous):

coughs dont worry about it ;)

OpenStudy (anonymous):

o.o Okie x.x

OpenStudy (anonymous):

im here to help not to be stood upon and rewarded when i shouldnt be

OpenStudy (anonymous):

i gave jakashaka 123 a meadeal

OpenStudy (anonymous):

medal

OpenStudy (anonymous):

thanks @hunter.glover

OpenStudy (anonymous):

how old are you

OpenStudy (anonymous):

16

OpenStudy (anonymous):

what is your name

OpenStudy (anonymous):

i am 13

OpenStudy (anonymous):

Stan is making a trail mix that has a combination of peanuts and raisins. The mix is 3 parts peanuts and 1 part raisins. How many cups of raisins does he need to use for a 16-cup recipe? 5cups 4 cups cups 64 cups

OpenStudy (anonymous):

my name is jacob and cool

OpenStudy (anonymous):

are you a girl or a boy

OpenStudy (anonymous):

boy

OpenStudy (anonymous):

Stan is making a trail mix that has a combination of peanuts and raisins. The mix is 3 parts peanuts and 1 part raisins. How many cups of raisins does he need to use for a 16-cup recipe? 5cups 4 cups cups 64 cups

OpenStudy (anonymous):

mixture is=peanuts:raisins=3:1 so in (3+1)cup i.e 4 cup there will be 1 cup raisin so for 16 cup recipe raisin required is=(1/4)*16=16/4=4 cup.

OpenStudy (anonymous):

4 i believe

OpenStudy (anonymous):

did he leave ?